# On the Convergence Rates of Federated Q-Learning across Heterogeneous Environments

### v1

@@ 6 Experiments
Description of the setup. In our experiments, we consider $K=5$ agents (Jin et al., 2022), each interacting with an independently and randomly generated $5\times 5$ maze environment $\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}}^{k},R,\gamma\rangle$ for $k\in\{1,2,\cdots,5\}$. The state set ${\mathcal{S}}$ contains $25$ cells that the agent is currently in. The action set contains 4 actions ${\mathcal{A}}=\{\text{left},\text{up},\text{right},\text{down}\}$. Thus, $|{\mathcal{S}}|\times|{\mathcal{A}}|=100$. We choose $\gamma=0.99$. For ease of verifying our theory, each entry of the reward $R\in{\mathbb{R}}^{100}$ is sampled from ${\rm Bern}(p=0.05)$, which slightly departs from a typical maze environment wherein only two state-action pairs have nonzero rewards. We choose this reward so that $\left\|{\Delta_{0}}\right\|_{\infty}\approx 100=\frac{1}{1-\gamma}$, which is the coarse upper bound of $\left\|{\Delta_{t}}\right\|_{\infty}$ for all $t$. For each agent $k$, its state transition probability vectors ${\mathcal{P}}^{k}$ are constructed on top of standard state transition probability vectors of maze environments incorporated with a drifting probability $0.1$ in each non-intentional action as in WindyCliff (Jin et al., 2022; Paul et al., 2019). In this way, the environment heterogeneity lies not only in the differences of the non-zero probability values (Jin et al., 2022; Paul et al., 2019) but also in the probability supports (i.e., the locations of non-zero entries). Our construction is more challenging: The environment heterogeneity $\kappa$ as per (2) of our environment construction was calculated to be $1.2$. Yet, the largest environment heterogeneity of the WindyCliff construction in Jin et al. (2022) is about 0.31.
@@
Two-phase phenomenon. We plot the evolutions of $\left\|{\Delta_{t}}\right\|_{\infty}$ for synchronous federated Q-learning under heterogeneous and homogeneous environments, respectively. Our results show that the sharp two-phase phenomenon mainly arises from environmental heterogeneity rather than sampling noise.
From Figure 2(a), it is clear that under the heterogeneous setting, for a given set of constant stepsizes $\lambda\in\{0.9,0.5,0.2,0.1,0.05\}$, the $\ell_{\infty}$-norm of $\Delta_{t}=Q^{*}-\bar{Q}_{t}$ decreases to a minimum point and then bounces back rapidly before stabilizing around some fixed errors. Moreover, we can see that different stepsizes give different minimum error. The smaller the stepsize, the smaller the minimum error; however, it takes longer to reach such minimum errors. In sharp contrast, as shown in Figure 2(b), there is no drastic bounce when the environments are homogeneous.
@@
Impacts of the synchronization period $E$. Furthermore, we test the impacts of the synchronization period $E$. As shown in Figure 3 and Figure 2(a), with $\lambda\in\{0.9,0.5,0.2,0.1,0.05\}$, as $E$ increases, the final error increases and saturates around 62 in the presence of environmental heterogeneity. For a homogeneous setting (results deferred to Appendix G.1), $E$ does not have a significant impact, which aligns with the observations in the existing literature on the homogeneous settings (Woo et al., 2023; Khodadadian et al., 2022).
Figure 3: Heterogeneous environments with varying $E$. From left to right, $E=1,20,40,$ and $\infty$ respectively
@@
As shown in Figures 2(a) and 3, in the presence of environmental heterogeneity, the smaller the stepsizes, the smaller error $\left\|{\Delta_{t}}\right\|_{\infty}$ can reach and less significant of the error bouncing in the second phase.
In our preliminary experiments, we tested small stepsizes $\lambda=1/T^{\alpha}$ for $\alpha\in\{0.4,0.5,\cdots,1\}$, which eventually lead to small errors yet at the cost of being extremely slow. Among these choices, $\lambda=1/\sqrt{T}$ has the fastest convergence performance yet is still $\approx 24$ up to iteration 20,000.
@@
Figure 4 illustrates two-phase training with different phase 1 stepsizes and phase 2 stepsize $\lambda=1/\sqrt{T}$ compared with using $\lambda=1/\sqrt{T}$ throughout. Overall, using $\lambda=1/\sqrt{T}$ throughout leads to the slowest convergence, highlighting the benefits of the two-phase training strategy. Among all two-phase stepsize choices, the stepsize of 0.05 in the first phase results in a longer phase 1 duration ($t_{0}=5550$ ) but the lowest final error (2.75327), suggesting a better convergence. We further test the convergence performance with respect to different target error levels and the convergence performance with different time-varying stepsizes, details can be found in Appendix G.2.
@@
Work RL
Algorithm
Hetero-
geneity
Optimality Lower bound Sampling Finite-time Task
Wang et al. (2023) TD(0) $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ Pred
Xie & Song (2023) PG $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ Pred, Plan
Zhang et al. (2024) SARSA $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ Pred, Plan
Khodadadian et al. (2022) TD,
Q-Learning
$\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ Pred, Plan
Jin et al. (2022) Q-Learning,
PG
$\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ Pred, Plan
Woo et al. (2023) Q-Learning $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ Pred, Plan
Zheng et al. (2023) Q-Learning $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ Pred, Plan
Our work Q-Learning $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Pred, Plan
Table 1: Comparison with related work. ‘Pred’ stands for prediction, and ‘PG’ stands for policy gradient.
@@ Appendix A Proof of Lemma 1
The update of $\Delta_{t+1}$ is as follows:
@@
$=\frac{1}{K}\sum_{k=1}^{K}(Q^{*}-((1-\lambda)Q_{t}^{k}+\lambda(R+\gamma\widetilde{P}_{t}^{k}Q_{t}^{k})))$
@@
$=(1-\lambda)\Delta_{t}+\gamma\lambda\frac{1}{K}\sum_{k=1}^{K}(\bar{P}V^{*}-\widetilde{P}_{t}^{k}V_{t}^{k})$
@@
recalling that $\Delta_{0}=Q^{*}-Q_{0},\text{ and }Q^{*}=R+\gamma\bar{P}V^{*}$.
@@ Appendix C Proof of Lemma 3
For general $t$, we have
@@
[…]
@@
For any state $s\in{\mathcal{S}}$, we have
@@
$=Q^{k}_{t}(s,a_{t}^{k}(s))-Q_{\chi(t)}^{k}(s,a_{\chi(t)}^{k}(s)),\text{ where }a_{t}^{k}(s)=\text{argmax}_{a^{\prime}}Q^{k}_{t}(s,a^{\prime})$
@@
=$\overset{(a)}{=}-\lambda Q^{k}_{t^{\prime}}(s,a_{t}^{k}(s))+\lambda\left(Q^{*}(s,a_{t}^{k}(s))-R(s,a_{t}^{k}(s))-\gamma\bar{P}(s,a_{t}^{k}(s))V^{*}+R(s,a_{t}^{k}(s))+\gamma\widetilde{P}_{t^{\prime}}^{k}(s,a_{t}^{k}(s))V_{t^{\prime}}^{k}\right)$
@@
where equality (a) follows from the Bellman equation  equation (3). Thus,
@@
Similarly, we have
@@
Plugging the bounds in equation (C) and in equation (C) back into equation (C), we get
@@ Appendix D Proof of Lemma 4
By invoking Hoeffding’s inequality, for any given $\delta\in(0,1)$, with probability at least $1-\delta$, it holds that
$\left\|{\gamma\lambda\sum_{j=\chi(i)}^{i-1}(1-\lambda)^{i-j-1}(\bar{P}-\widetilde{P}_{j}^{k})V^{*}}\right\|_{\infty}$ $\leq\frac{\gamma}{1-\gamma}\lambda\sum_{j=\chi(i)}^{i-1}(1-\lambda)^{i-1-j}\kappa+\frac{\gamma}{1-\gamma}\sqrt{\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}$
@@
for all $(s,a)\in{\mathcal{S}}\times{\mathcal{A}},i\in[T],k\in[K]$. In addition, we have
@@
Combining the bounds in equation (18), equation (D), and equation (20), we get
@@ Appendix E Proof of Theorem 1
$\left\|{\Delta_{t+1}}\right\|_{\infty}$ $\leq(1-\lambda)^{t+1}\left\|{\Delta_{0}}\right\|_{\infty}+\left\|{\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\frac{1}{K}\sum_{k=1}^{K}(\bar{P}-\widetilde{P}_{t-i}^{k})V^{*}}\right\|_{\infty}$
$\quad+\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{t-i}^{k}(V^{*}-V_{t-i}^{k})}\right\|_{\infty}.$
@@
Since $0\leq Q_{0}(s,a)\leq\frac{1}{1-\gamma}$, the first term can be bounded as
$(1-\lambda)^{t+1}\left\|{\Delta_{0}}\right\|_{\infty}\leq(1-\lambda)^{t+1}\frac{1}{1-\gamma}.$ (22)
To bound the second term $\left\|{\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\frac{1}{K}\sum_{k=1}^{K}(\bar{P}-\widetilde{P}_{t-i}^{k})V^{*}}\right\|_{\infty}$, we have
@@
$\left\|{\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\frac{1}{K}\sum_{k=1}^{K}(\bar{P}-\widetilde{P}_{t-i}^{k})V^{*}}\right\|_{\infty}$ $\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}.$ (23)
To bound the third term $\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{t-i}^{k}(V^{*}-V_{t-i}^{k})}\right\|_{\infty}$, following the roadmap of Woo et al. (2023), we divide the summation into two parts as follows. For any $\beta E\leq t\leq T$, we have
$\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{t-i}^{k}(V^{*}-V_{t-i}^{k})}\right\|_{\infty}$
@@
Thus, with probability at least $(1-2\delta)$,
@@
The third term can be bounded as
$\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{i}^{k}(V^{*}-V_{i}^{k})}\right\|_{\infty}$
@@
Combing the bounds for terms (22), (23), and (E), we get the following recursion holds for all rounds $T$ with probability at least $(1-3\delta)$:
$\left\|{\Delta_{t+1}}\right\|_{\infty}$ $\leq(1-\lambda)^{t+1}\frac{1}{1-\gamma}+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}+\frac{\gamma}{1-\gamma}(1-\lambda)^{t-\chi(t)+\beta E}$
@@
$\quad+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}.$
@@
$+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}.$ (26)
@@
$\quad+\left.\left(\frac{6\gamma^{2}\lambda}{1-\gamma}\sqrt{E-1}+\frac{\gamma^{2}\sqrt{\lambda}}{1-\gamma}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}\right)$
$\leq\frac{1}{1-\gamma}(\frac{1+\gamma}{2})^{\sqrt{\frac{\lambda T}{1-\gamma}}}+\frac{4}{(1-\gamma)^{2}}(1-\lambda)^{\sqrt{\frac{(1-\gamma)T}{2\lambda}}}+\frac{2\gamma^{2}}{(1-\gamma)^{2}}(6\lambda^{2}(E-1)^{2}+\lambda(E-1))\kappa$
$\quad+\left(\frac{12\gamma^{2}\lambda}{(1-\gamma)^{2}}\sqrt{E-1}+\frac{2\gamma^{2}\sqrt{\lambda}}{(1-\gamma)^{2}}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}+\frac{2\gamma}{(1-\gamma)^{2}}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}$
$\leq\frac{1}{1-\gamma}\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}+\frac{4}{(1-\gamma)^{2}}\exp\left\{-\sqrt{(1-\gamma)\lambda T}\right\}$
@@
$\quad+\left(\frac{12\gamma^{2}\lambda}{(1-\gamma)^{2}}\sqrt{E-1}+\frac{2\gamma^{2}\sqrt{\lambda}}{(1-\gamma)^{2}}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}+\frac{2\gamma}{(1-\gamma)^{2}}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}$
@@
$\quad+\left(\frac{12\gamma^{2}\lambda}{(1-\gamma)^{2}}\sqrt{E-1}+\frac{2\gamma^{2}\sqrt{\lambda}}{(1-\gamma)^{2}}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}+\frac{2\gamma}{(1-\gamma)^{2}}\sqrt{\frac{1}{K}\lambda\log\frac{|S||A|TK}{\delta}}.$

### v2

@@ 6 Experiments
Description of the setup. In our experiments, we consider $K=20$ agents (Jin et al., 2022), each interacting with an independently and randomly generated $5\times 5$ maze environment $\langle{\mathcal{S}},{\mathcal{A}},{\mathcal{P}}^{k},R,\gamma\rangle$ for $k\in\{1,2,\cdots,20\}$. The state set $\mathcal{S}$} contains $25$ cells that the agent is currently in. The action set contains 4 actions ${\mathcal{A}}=\{\text{left},\text{up},\text{right},\text{down}\}$. Thus, $|{\mathcal{S}}|\times|{\mathcal{A}}|=100$. We choose $\gamma=0.99$. For ease of verifying our theory, each entry of the reward $R\in{\mathbb{R}}^{100}$ is sampled from ${\rm Bern}(p=0.05)$, which slightly departs from a typical maze environment wherein only two state-action pairs have nonzero rewards. We choose this reward so that $\left\|{\Delta_{0}}\right\|_{\infty}\approx 100=\frac{1}{1-\gamma}$, which is the coarse upper bound of $\left\|{\Delta_{t}}\right\|_{\infty}$ for all $t$. For each agent $k$, its state transition probability vectors ${\mathcal{P}}^{k}$ are constructed on top of standard state transition probability vectors of maze environments incorporated with a drifting probability $0.1$ in each non-intentional action as in WindyCliff (Jin et al., 2022; Paul et al., 2019). In this way, the environment heterogeneity lies not only in the differences of the non-zero probability values (Jin et al., 2022; Paul et al., 2019) but also in the probability supports (i.e., the locations of non-zero entries). Our construction is more challenging: The environment heterogeneity $\kappa$ as per (2) of our environment construction was calculated to be $1.2$. Yet, the largest environment heterogeneity of the WindyCliff construction in Jin et al. (2022) is about 0.31.
@@
Convergence behavior and two-phase phenomenon.
We demonstrate through numerical simulations that our analysis aligns with the observed behaviors. For algorithms with a time-invariant stepsize, convergence requires sufficiently small stepsizes and a sufficiently large number of iterations $T$.
To explore the impact of stepsizes on convergence, we use $\lambda\in\{0.9,0.5,0.2,0.1,0.05\}$, spanning a range within $(0,1)$. As shown in Figure 2(a), these stepsizes are not sufficiently small, leading to a two-phase phenomenon: the $\ell_{\infty}$-norm of $\Delta_{t}=Q^{*}-\bar{Q}_{t}$ has a rapid decay in the first phase followed by a bounce back in the second phase. This phenomenon is distinctive to heterogeneous settings. In contrast, Figure 2(b) indicates that in homogeneous environments, no drastic bounce occurs, irrespective of the stepsize. Note that if multiple plots on the same page share the same legend, we display the legend only once for clarity.
Figure 4(a) (light blue curve) demonstrates that with a sufficiently small stepsize, such as $\lambda=\frac{1}{\sqrt{T}}$, the error continuously decreases, reaching approximately 24 by iteration 20,000.
@@
Impacts of the synchronization period $E$.
In homogeneous settings (refer to Figure 5 in Appendix G.1), the synchronization period $E$ has negligible impact, consistent with prior findings in the literature (Woo et al., 2023; Khodadadian et al., 2022). However, under heterogeneous conditions, larger $E$ values lead to increased final error across the five constant stepsizes, as depicted in Figure 3 and Figure 2(a). This degradation persists even with time-decaying stepsizes $\lambda_{t}=\frac{1}{\sqrt{t+1}}$, as shown in Figure 6. We hypothesize that larger $E$ values require either smaller or more rapidly decaying stepsizes to mitigate the degradation caused by increased synchronization periods.
@@
As shown in Figures 2(a) and 3, in the presence of environmental heterogeneity, the smaller the stepsizes, the smaller error $\left\|{\Delta_{t}}\right\|_{\infty}$ can reach and less significant of the error bouncing in the second phase. In our preliminary experiments, we tested small stepsizes $\lambda=1/T^{\alpha}$ for $\alpha\in\{0.4,0.5,\cdots,1\}$, which eventually lead to small errors yet at the cost of being extremely slow. Among these choices, $\lambda=1/\sqrt{T}$ has the fastest convergence performance yet is still $\approx 24$ up to iteration 20,000.
(a) E=1
(b) E=20
(c) E=40
[…]
Figure 3: Convergence behavior for constant stepsizes $(0.05,0.1,0.2,0.5,0.9)$ under various synchronization intervals $E$ $(1,20,40,\infty)$. In heterogeneous settings, higher $E$ and larger $\lambda$ lead to higher residual errors.
@@
Figure 4(a) illustrates two-phase training with different phase 1 stepsizes and phase 2 stepsize $\lambda=1/\sqrt{T}$ compared with using $\lambda=1/\sqrt{T}$ throughout. Overall, using $\lambda=1/\sqrt{T}$ throughout leads to the slowest convergence, highlighting the benefits of the two-phase training strategy. Among all two-phase stepsize choices, the stepsize of 0.05 in the first phase results in a longer phase 1 duration ($t_{0}=5550$) but the lowest final error (2.75327), suggesting a better convergence. We further test the convergence performance with respect to different target error levels, details can be found in Appendix G.3.
We also evaluated the two-phase training strategy using various time-decaying step sizes, including $\frac{1}{\sqrt{t+1}}$, $\frac{c+1}{t+c}$, $\frac{1}{t+1}$, and $\frac{1}{(t+1)^{0.7}}$. In all cases, Figure 4 shows the two-phase training has an advantage.
@@
Acknowledgments
We thank He Cheng for his valuable assistance on the experiments. P. Yang is supported in part by the National Key R&D Program of China 2024YFA1015800, and Tsinghua University Dushi Program 2025Z11DSZ001. L. Su is supported in part by NSF CAREER CIF-2340482.
@@ Appendix A Proof of Lemma 1
The evolution of $\Delta_{t+1}$ can be decomposed as follows:
@@
$\overset{(a)}{=}\frac{1}{K}\sum_{k=1}^{K}(Q^{*}-((1-\lambda)Q_{t}^{k}+\lambda(R+\gamma\widetilde{P}_{t}^{k}V_{t}^{k})))$
@@
$\overset{(b)}{=}(1-\lambda)\Delta_{t}+\gamma\lambda\frac{1}{K}\sum_{k=1}^{K}(\bar{P}V^{*}-\widetilde{P}_{t}^{k}V_{t}^{k})$
@@
where equality (a) follows from Eq. (4), equality (b) follows from the Bellman optimality equation in Eq. (3), and the last equality follows from unrolling the updates $t+1$ times and using the fact that $\Delta_{0}=Q^{*}-Q_{0}$.
@@ Appendix C Proof of Lemma 3
For general $t$, we anchor the error term to that of the synchronization round as follows:
@@
$\overset{(a)}{\leq}\left\|{\Delta_{\chi(t)}}\right\|_{\infty}+\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{t}^{k}(V_{\chi(t)}^{k}-V^{k}_{t})}\right\|_{\infty}$
@@
where inequality (a) follows from Eq. (C). For any state $s\in{\mathcal{S}}$, let $a_{t}^{k}(s)\in\arg\max_{a\in{\mathcal{A}}}Q_{t}^{k}(s,a)$ for all $t$ and $k$. We have
@@
$=Q^{k}_{t}(s,a_{t}^{k}(s))-Q_{\chi(t)}^{k}(s,a_{\chi(t)}^{k}(s))$
@@
$\overset{(a)}{=}-\lambda Q^{k}_{t^{\prime}}(s,a_{t}^{k}(s))+\lambda\left(Q^{*}(s,a_{t}^{k}(s))-R(s,a_{t}^{k}(s))-\gamma\bar{P}(s,a_{t}^{k}(s))V^{*}+R(s,a_{t}^{k}(s))+\gamma\widetilde{P}_{t^{\prime}}^{k}(s,a_{t}^{k}(s))V_{t^{\prime}}^{k}\right)$
$=\lambda\Delta_{t^{\prime}}^{k}(s,a_{t}^{k}(s))+\gamma\lambda\left(\widetilde{P}_{t^{\prime}}^{k}(s,a_{t}^{k}(s))V_{t^{\prime}}^{k}-\bar{P}(s,a_{t}^{k}(s))V^{*}\right)$
@@
where equality (a) follows from the Bellman equation Eq. (3), and the last inequality follows from Eq.(C) and that
$\gamma\lambda\widetilde{P}_{t^{\prime}}^{k}(s,a_{t}^{k}(s))(V_{t^{\prime}}^{k}-V^{*})\leq\gamma\lambda\|\widetilde{P}_{t^{\prime}}^{k}(s,a_{t}^{k}(s))\|_{1}\|V_{t^{\prime}}^{k}-V^{*}\|_{\infty}\leq\gamma\|V_{t^{\prime}}^{k}-V^{*}\|_{\infty}.$
Thus, $V^{k}_{t}(s)-V_{\chi(t)}^{k}(s)$ can be upper bounded as
@@
Similarly, we have the following lower bound:
@@
Plugging the bounds in Eq. (C) and in Eq. (C) back into Eq. (C), we get
@@
$\quad+\gamma\lambda\frac{1}{K}\sum_{k=1}^{K}\max_{s,a}{\left|\sum_{t^{\prime}=\chi(t)}^{t-1}\left(\widetilde{P}_{t^{\prime}}^{k}(s,a)-\bar{P}(s,a)\right)V^{*}\right|},$
proving the lemma.
@@ Appendix D Proof of Lemma 4
For the second term, we have
$\left\|{\gamma\lambda\sum_{j=\chi(i)}^{i-1}(1-\lambda)^{i-j-1}(\bar{P}-\widetilde{P}_{j}^{k})V^{*}}\right\|_{\infty}$ (20)
$\leq$ $\left\|{\gamma\lambda\sum_{j=\chi(i)}^{i-1}(1-\lambda)^{i-j-1}(\bar{P}-{P}_{j}^{k})V^{*}}\right\|_{\infty}+\left\|{\gamma\lambda\sum_{j=\chi(i)}^{i-1}(1-\lambda)^{i-j-1}({P}_{j}^{k}-\widetilde{P}_{j}^{k})V^{*}}\right\|_{\infty}$ (21)
$\leq$ $\frac{\gamma}{1-\gamma}\lambda\sum_{j=\chi(i)}^{i-1}(1-\lambda)^{i-1-j}\kappa+\frac{\gamma}{1-\gamma}\sqrt{\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}$ (22)
@@
for all $(s,a)\in{\mathcal{S}}\times{\mathcal{A}},i\in[T],k\in[K]$. From Eq. (21) to Eq. (22), since for each timestep, each agent independently samples the next state for all state-action pairs to form the sample transition matrix, we can use Hoeffding’s inequality by treating $\lambda(1-\lambda)^{i-j-1}({P}_{j}^{k}-\widetilde{P}_{j}^{k})V^{*}$ as the independent random variables with their absolute values bounded by $\lambda(1-\lambda)^{i-j-1}\left\|{V^{*}}\right\|_{\infty}$.
In addition, we have
@@
Combining the bounds in Eq. (19), Eq. (20), and Eq. (24), we get
@@ Appendix E Proof of Theorem 1
$\left\|{\Delta_{t+1}}\right\|_{\infty}$ $\leq\underbrace{(1-\lambda)^{t+1}\left\|{\Delta_{0}}\right\|_{\infty}}_{\textnormal{(I.1)}}+\underbrace{\left\|{\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\frac{1}{K}\sum_{k=1}^{K}(\bar{P}-\widetilde{P}_{t-i}^{k})V^{*}}\right\|_{\infty}}_{\textnormal{(I.2)}}$
$\quad+\underbrace{\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{t-i}^{k}(V^{*}-V_{t-i}^{k})}\right\|_{\infty}}_{\textnormal{(I.3)}}.$ (26)
@@
Bounding (I.1). Since $0\leq Q_{0}(s,a)\leq\frac{1}{1-\gamma}$, the first term can be bounded as
$\textnormal{(I.1)}=(1-\lambda)^{t+1}\left\|{\Delta_{0}}\right\|_{\infty}\leq(1-\lambda)^{t+1}\frac{1}{1-\gamma}.$ (27)
Bounding (I.2). To bound the second term (I.2) in Eq. (E), we have
@@
$\textnormal{(I.2)}=\left\|{\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\frac{1}{K}\sum_{k=1}^{K}(\bar{P}-\widetilde{P}_{t-i}^{k})V^{*}}\right\|_{\infty}$ $\leq\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}.$ (28)
Bounding (I.3). To bound the third term (I.3) in Eq. (E), following the roadmap of Woo et al. (2023), we divide the summation into two parts as follows. For any $\beta E\leq t\leq T$, we have
$\textnormal{(I.3)}=\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{t-i}^{k}(V^{*}-V_{t-i}^{k})}\right\|_{\infty}$
@@
Thus, by applying the union bound, we get with probability at least $(1-2\delta)$,
@@
Hence, the third term (I.3) in Eq. (E) can be bounded as
$\textnormal{(I.3)}=\sum_{i=0}^{t}(1-\lambda)^{i}\lambda\gamma\left\|{\frac{1}{K}\sum_{k=1}^{K}\widetilde{P}_{i}^{k}(V^{*}-V_{i}^{k})}\right\|_{\infty}$
@@
Combing the bounds of (I.1), (I.2), and (I.3) in Eq. (E). Combing the bounds for terms (27), (28), and (E), we get the following recursion holds for all rounds $T$ with probability at least $(1-3\delta)$:
$\left\|{\Delta_{t+1}}\right\|_{\infty}$ $\leq(1-\lambda)^{t+1}\frac{1}{1-\gamma}+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}+\frac{\gamma}{1-\gamma}(1-\lambda)^{t-\chi(t)+\beta E}$
@@
$\quad+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}.$
@@
$+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}.$ (31)
@@
$\quad+\left.\left(\frac{6\gamma^{2}\lambda}{1-\gamma}\sqrt{E-1}+\frac{\gamma^{2}\sqrt{\lambda}}{1-\gamma}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}+\frac{\gamma}{1-\gamma}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}\right)$
$\leq\frac{1}{1-\gamma}\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}+\frac{4}{(1-\gamma)^{2}}\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}$
@@
$\quad+\left(\frac{12\gamma^{2}\lambda}{(1-\gamma)^{2}}\sqrt{E-1}+\frac{2\gamma^{2}\sqrt{\lambda}}{(1-\gamma)^{2}}\right)\sqrt{\lambda(E-1)\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}+\frac{2\gamma}{(1-\gamma)^{2}}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}}$
@@
$\quad{+\left(\frac{14\gamma^{2}\lambda}{(1-\gamma)^{2}}\sqrt{E-1}\right)\sqrt{\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}}+\frac{2\gamma}{(1-\gamma)^{2}}\sqrt{\frac{1}{K}\lambda\log\frac{|{\mathcal{S}}||{\mathcal{A}}|TK}{\delta}},$
where the second inequality follows from
$(\frac{1+\gamma}{2})^{\sqrt{\frac{\lambda T}{1-\gamma}}}=(1-\frac{1-\gamma}{2})^{\sqrt{\frac{\lambda T}{1-\gamma}}}\leq\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\},$
$(1-\lambda)^{\beta E}\leq\exp\left\{-\lambda\sqrt{\frac{(1-\gamma)T}{2\lambda}}\right\}\leq\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}.$
By the assumption that $(E-1)\leq\frac{1}{K\lambda}$, the above can be further simplified as
$\left\|{\Delta_{T}}\right\|_{\infty}$ $\leq\frac{4}{(1-\gamma)^{2}}\exp\left\{-\frac{1}{2}\sqrt{(1-\gamma)\lambda T}\right\}+\frac{14\gamma^{2}}{(1-\gamma)^{2}}\lambda(E-1)\kappa+\frac{16}{(1-\gamma)^{2}}\sqrt{\frac{\lambda}{K}\log\frac{|{\mathcal{S}}||{\mathcal{A}}|KT}{\delta}}.$